# cvprbluergb0.21,0.49,0.74

### v3

@@ 3.1.3 Heterogeneous Robot Platforms
Liu et al1 0.307 0.095 0.629 0.122 0.235 0.260 0.889 0.757 0.978 1.454 0.401 0.934 0.693
Weitong et al2 0.26 0.096 0.617 0.120 0.222 0.402 0.998 0.770 1.586 1.254 0.577 1.056 0.789
Kim et al3 0.331 0.092 0.787 0.123 0.270 0.279 10.628 22.425 7.147 2.100 0.650 1.068 4.996
Yibin et al4 1.060 0.220 0.750 0.470 0.620 9.140 4.920 0.280 24.460 2.990 5.500 1.340 5.096
Zhong et al5 1.205 0.695 - 1.175 1.72 2.08 0.889 0.778 1.13 - - - 1.154*
@@
Peng et al1 1.063 1.637 0.503 0.44 0.153 0.827 0.038 0.195 0.070 0.436
Thien et al2 1.081 2.054 1.733 1.054 10.532 7.692 0.753 1.228 1.209 2.543
Jiang et al3 1.019 1.126 1.911 2.341 3.757 11.821 2.154 0.604 4.010 2.956
Li et al4 5.768 7.834 1.757 1.295 5.370 10.766 - 30.07 - 17.76*
@@ 4.1 Accuracy Evaluation
Liu et al1 $R_{p}\uparrow$ 0.811 0.865 0.736 0.747 0.504 0.816 0.157 0.135 0.135 0.396 0.529 0.371 0.440
Weitong et al2 0.773 0.838 0.690 0.871 0.739 0.152 0.153 0.149 0.081 0.407 0.456 0.345 0.405
Kim et al4 0.747 0.865 0.737 0.870 0.689 0.777 0.014 0.001 0.010 0.410 0.285 0.345 0.379
Yibin et al3 0.650 0.783 0.557 0.721 0.490 0.481 0.219 0.560 0.019 0.175 0.260 0.370 0.382
Zhong et al5 0.567 0.683 0.204 0.680 0.426 0.447 0.103 0.084 0.122 0 0 0 0.201
Liu et al1 $R_{r}\uparrow$ 0.888 0.893 0.816 0.857 0.778 0.861 0.692 0.719 0.652 0.753 0.689 0.643 0.743
Weitong et al2 0.886 0.892 0.811 0.893 0.840 0.598 0.688 0.720 0.514 0.766 0.706 0.644 0.722
Kim et al4 0.680 0.893 0.816 0.893 0.803 0.808 0.198 0.125 0.300 0.802 0.422 0.642 0.548
Yibin et al3 0.624 0.707 0.462 0.733 0.540 0.467 0.862 0.962 0.496 0.477 0.342 0.213 0.568
Zhong et al5 0.731 0.745 0.437 0.757 0.603 0.523 0.688 0.704 0.643 0 0 0 0.437
@@
Peng et al1 $R_{p}\uparrow$ 0.357 0.227 0.264 0.203 0.536 0.270 0.699 0.893 0.806 0.555
Thien et al3 0.045 0.070 0.240 0.156 0.131 0.075 0 0.031 0 0.066
Jiang et al 4 0.046 0.039 0.194 0.088 0.242 0.095 0 0 0 0.0585
Li et al2 0.342 0.187 0.257 0.208 0.322 0.142 0 0.006 0 0.123
Peng et al1 $R_{r}\uparrow$ 0.641 0.581 0.744 0.650 0.878 0.670 0.975 0.975 0.974 0.836
Thien et al3 0.413 0.445 0.610 0.269 0.474 0.487 0.177 0.315 0 0.276
Jiang et al4 0.453 0.452 0.619 0.252 0.542 0.577 0.002 0 0.157 0.242
Li et al2 0.642 0.574 0.657 0.651 0.773 0.660 0.0 0.305 0 0.381
@@ Appendix C Derivation of Robustness Metric
Liu et al1 ATE 0.307 0.095 0.629 0.122 0.235 0.260 0.889 0.757 0.978 1.454 0.401 0.934 0.693
Yibin et al4 1.060 0.220 0.750 0.470 0.620 9.140 4.920 0.280 24.460 2.990 5.500 1.340 5.096
Weitong et al2 0.26 0.096 0.617 0.120 0.222 0.402 0.998 0.770 1.586 1.254 0.577 1.056 0.789
Kim et al3 0.331 0.092 0.787 0.123 0.270 0.279 10.628 22.425 7.147 2.100 0.650 1.068 4.996
Zhong et al5 1.205 0.695 - 1.175 1.72 2.08 0.889 0.778 1.13 - - - 1.154*
Liu et al1 RPE 0.038 0.032 0.055 0.028 0.048 0.040 0.191 0.174 0.188 0.088 0.059 0.148 0.322
Yibin et al3 0.130 0.090 0.150 0.130 0.200 0.200 0.040 0.040 0.160 0.630 0.280 0.180 0.593
Weitong et al2 0.038 0.032 0.056 0.029 0.049 0.046 0.190 0.183 0.243 0.086 0.054 0.166 0.349
Kim et al4 0.098 0.032 0.055 0.028 0.054 0.046 0.861 0.535 0.401 0.093 0.26 0.167 0.824
Zhong et al5 0.157 0.062 - 0.079 0.1062 0.0937 0.706 0.691 0.617 - - - 0.770*
@@
Peng et al1 ATE 1.063 1.637 0.503 0.44 0.153 0.827 0.038 0.195 0.070 0.436
Jiang et al3 1.019 1.126 1.911 2.341 3.757 11.821 2.154 0.604 4.010 2.956
Thien et al2 1.081 2.054 1.733 1.054 10.532 7.692 0.753 1.228 1.209 2.543
Li et al4 5.768 7.834 1.757 1.295 5.370 10.766 - 30.07 - 17.76*
Peng et al1 RPE 0.058 0.063 0.051 0.149 0.026 0.064 0.002 0.014 0.01 0.039
Jiang et al3 0.190 0.203 0.258 0.307 0.884 3.427 3.982 0.792 7.477 2.859
Thien et al2 0.197 0.186 0.181 0.231 0.071 0.279 0.471 0.007 0.777 0.296
Li et al4 0.088 0.088 0.124 0.160 0.911 0.478 - 0.347 - 0.328*

### v4

@@ 3.1.3 Heterogeneous Robot Platforms
Liu et al1 0.307 0.095 0.629 0.122 0.235 0.260 0.889 0.757 0.978 1.454 0.401 0.934 0.588
Weitong et al2 0.26 0.096 0.617 0.120 0.222 0.402 0.998 0.770 1.586 1.254 0.577 1.056 0.663
Kim et al3 0.331 0.092 0.787 0.123 0.270 0.279 10.628 22.425 7.147 2.100 0.650 1.068 3.825
Yibin et al4 1.060 0.220 0.750 0.470 0.620 9.140 4.920 0.280 24.460 2.990 5.500 1.340 4.312
Zhong et al5 1.205 0.695 - 1.175 1.72 2.08 0.889 0.778 1.13 - - - 1.209*
@@
Peng et al1 1.063 1.637 0.503 0.44 0.153 0.827 0.038 0.195 0.070 0.547
Thien et al2 1.081 2.054 1.733 1.054 10.532 7.692 0.753 1.228 1.209 3.037
Jiang et al3 1.019 1.126 1.911 2.341 3.757 11.821 2.154 0.604 4.010 3.193
Li et al4 5.768 7.834 1.757 1.295 5.370 10.766 - 30.07 - 8.98*
@@ 4.1 Accuracy Evaluation
Liu et al1 $R_{p}\uparrow$ 0.811 0.865 0.736 0.747 0.504 0.816 0.157 0.135 0.135 0.396 0.529 0.371 0.516
Weitong et al2 0.773 0.838 0.690 0.871 0.739 0.152 0.153 0.149 0.081 0.407 0.456 0.345 0.471
Kim et al4 0.747 0.865 0.737 0.870 0.689 0.777 0.014 0.001 0.010 0.410 0.285 0.345 0.479
Yibin et al3 0.650 0.783 0.557 0.721 0.490 0.481 0.219 0.560 0.019 0.175 0.260 0.370 0.440
Zhong et al5 0.567 0.683 0.204 0.680 0.426 0.447 0.103 0.084 0.122 0 0 0 0.276
Liu et al1 $R_{r}\uparrow$ 0.888 0.893 0.816 0.857 0.778 0.861 0.692 0.719 0.652 0.753 0.689 0.643 0.770
Weitong et al2 0.886 0.892 0.811 0.893 0.840 0.598 0.688 0.720 0.514 0.766 0.706 0.644 0.746
Kim et al4 0.680 0.893 0.816 0.893 0.803 0.808 0.198 0.125 0.300 0.802 0.422 0.642 0.615
Yibin et al3 0.624 0.707 0.462 0.733 0.540 0.467 0.862 0.962 0.496 0.477 0.342 0.213 0.573
Zhong et al5 0.731 0.745 0.437 0.757 0.603 0.523 0.688 0.704 0.643 0 0 0 0.485
@@
Peng et al1 $R_{p}\uparrow$ 0.357 0.227 0.264 0.203 0.536 0.270 0.699 0.893 0.806 0.472
Thien et al3 0.045 0.070 0.240 0.156 0.131 0.075 0 0.031 0 0.083
Jiang et al 4 0.046 0.039 0.194 0.088 0.242 0.095 0 0 0 0.078
Li et al2 0.342 0.187 0.257 0.208 0.322 0.142 0 0.006 0 0.162
Peng et al1 $R_{r}\uparrow$ 0.641 0.581 0.744 0.650 0.878 0.670 0.975 0.975 0.974 0.787
Thien et al3 0.413 0.445 0.610 0.269 0.474 0.487 0.177 0.315 0 0.354
Jiang et al4 0.453 0.452 0.619 0.252 0.542 0.577 0.002 0 0.157 0.339
Li et al2 0.642 0.574 0.657 0.651 0.773 0.660 0.0 0.305 0 0.473
@@ Appendix C Derivation of Robustness Metric
Liu et al1 ATE 0.307 0.095 0.629 0.122 0.235 0.260 0.889 0.757 0.978 1.454 0.401 0.934 0.588
Yibin et al4 1.060 0.220 0.750 0.470 0.620 9.140 4.920 0.280 24.460 2.990 5.500 1.340 4.312
Weitong et al2 0.26 0.096 0.617 0.120 0.222 0.402 0.998 0.770 1.586 1.254 0.577 1.056 0.663
Kim et al3 0.331 0.092 0.787 0.123 0.270 0.279 10.628 22.425 7.147 2.100 0.650 1.068 3.825
Zhong et al5 1.205 0.695 - 1.175 1.72 2.08 0.889 0.778 1.13 - - - 1.209*
Liu et al1 RPE 0.038 0.032 0.055 0.028 0.048 0.040 0.191 0.174 0.188 0.088 0.059 0.148 0.091
Yibin et al3 0.130 0.090 0.150 0.130 0.200 0.200 0.040 0.040 0.160 0.630 0.280 0.180 0.186
Weitong et al2 0.038 0.032 0.056 0.029 0.049 0.046 0.190 0.183 0.243 0.086 0.054 0.166 0.097
Kim et al4 0.098 0.032 0.055 0.028 0.054 0.046 0.861 0.535 0.401 0.093 0.26 0.167 0.219
Zhong et al5 0.157 0.062 - 0.079 0.1062 0.0937 0.706 0.691 0.617 - - - 0.313*
@@
Peng et al1 ATE 1.063 1.637 0.503 0.44 0.153 0.827 0.038 0.195 0.070 0.547
Jiang et al3 1.019 1.126 1.911 2.341 3.757 11.821 2.154 0.604 4.010 3.193
Thien et al2 1.081 2.054 1.733 1.054 10.532 7.692 0.753 1.228 1.209 3.037
Li et al4 5.768 7.834 1.757 1.295 5.370 10.766 - 30.07 - 8.98*
Peng et al1 RPE 0.058 0.063 0.051 0.149 0.026 0.064 0.002 0.014 0.01 0.048
Jiang et al3 0.190 0.203 0.258 0.307 0.884 3.427 3.982 0.792 7.477 1.947
Thien et al2 0.197 0.186 0.181 0.231 0.071 0.279 0.471 0.007 0.777 0.266
Li et al4 0.088 0.088 0.124 0.160 0.911 0.478 - 0.347 - 0.314*